A car manufacturer produces three models in equal quantities. One model has an economy rating of 17 miles per gallon, and the second model is rated for 19 mpg. If government regulations require the manufacturer to have a fleet average that exceeds 21 mpg, what economy rating is required for the third model?
Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

(17 + 19 + t)/3 > 21
17 + 19 + t > 63
t > 63 - 36
---
t > 27 mpg
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by stanbon(75887)

A car manufacturer produces three models in equal quantities. One model has an economy rating of 17 miles per gallon, and the second model is rated for 19 mpg. If government regulations require the manufacturer to have a fleet average that exceeds 21 mpg, what economy rating is required for the third model?
-------------
Let the 3rd model rating be "x":
------
Equation:
(17 + 19 + x)/3 = 21
------
36 + x = 3*21
x = 63-36
x = 27 mph
-------------------
Cheers,
Stan H.
================